From GlueXWiki
Jump to: navigation, search

November 5, 2015 Drift Chamber meeting


  1. Instructions for Bluejeans meeting connection
  2. Meeting ID: 290664653
  3. To join via a Web Browser, go to the page [1] https://bluejeans.com/290664653.


  1. CDC update (Mike, Simon)
  2. FDC update x-resolution improvement (Lubomir)
  3. Electronics (Fernando, Chris, Nick)
  4. Engineering (Dave)
  5. Other


Participants: Curtis, Naomi, Mike (CMU), Sean (NU), Fernando, Cody, Marc, Dave, Chris, Nick, Simon, Beni, and Lubomir (JLab).

CDC update

- Mike: applying deformation corrections improves the CDC resolution to its design value of ~150 microns. This is demonstrated in his talk at the DNP meeting: last two slides where he first applies the corrections and then the alignment offsets, resulting in 193 down to 134 microns resolution.

- Mike send a list of not working channels to Nick. Nick fixed some but some will require access to the upstream side of the CDC. This is not possible now because the target is in. By Monday will decide if we need to move the target back (will require re-surveying). In the mean time we will postpone the installation of the spare FDC package on the platform, to make sure there's space if needed.

- Naomi asked how to prepare the CDC configuration file for the new firmware. Beni: it's up to Sasha, but in the mean time she can send the configuration parameters to Sergey to be hard coded.

FDC update

- There's an improvement in the Cathode X-resolution (cathode reconstructed - nominal wire position). See plot above - blue is old result, red is the new one; the two distributions have the same normalization. This is a result of systematic corrections vs distance to the middle of the strip (as discussed in the Collaboration meeting) and some additional channel calibrations.

- Simon and Lubomir are working on implementing Lubomir's code for the 3d-point reconstruction into the Simon's tracking. First we will compare the tracking results of the two codes using straight tracks.

- Today at the run coordination meeting we decided to move the spare FDC package to the Hall on Monday, but it will be postponed. We need to install 3 fADC125 units in some crate and we discussed the options - most likely in the start counter crate. We need one additional HV module to be installed somewhere, and 4 extra LV channels. This will be decided with Fernando, Nick next week.


- Fernando looked with a scope on the CDC and FDC pre-amp response function using a pulser. There's no difference between the two - so the tails must come from the detector. More detailed studies will be done in the tagger where we have a DAQ working, so we will include the fADCs into the chain. Nick and Lubomir will work on this tomorrow.

- Cody uploaded the latest firmware into the CDC crates. He corrected a problem found by Naomi in case of multiple peak processing. Cody also implemented busy signal as trigger hold-off, but it requires update of the TI firmware. If this doesn't work, a fixed 5 msec hold-off can be used. All the detectors must be ready for the run in three weeks from now (as decided on the run coordination meeting today) and this includes the new firmware.


- Dave: the NBX serial issue is still there, but not visible now after a temporary fix.


- Beni brought the subject of instrumentation papers, one for each CDC and FDC. Fernando: is it worth writing separate paper on the electronics? Most likely it is, since the only "publication" so far is a poster at the IAAA conference by Gerard years ago.